OpenStudy (anonymous):

Solve the equation for y if |y + 3| = 9 y = 6, -12 y = 6, -6 y = 12, -12 y = 6

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

i need this bad.

OpenStudy (mathmate):

\(|y+3|=9\) means y+3=9 \(or\) -(y+3)=9 Solve each case and choose the correct answer.

OpenStudy (anonymous):

12,-12

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Hmmm... Let me see: if y=12, then \(12+3\ne 9 \), so 12 does not work. now if y=-12, then \(-12+3= 9 \) yay, that's one of them!